Overhead Variances, Four-Variance Analysis, Journal Entries

Laughlin, Inc., uses a standard costing system. The predetermined overhead rates are calculated using practical capacity. Practical capacity for a year is defined as 1,000,000 units requiring 200,000 standard direct labor hours. Budgeted overhead for the year is $750,000, of which $300,000 is fixed overhead. During the year, 900,000 units were produced using 190,000 direct labor hours. Actual annual overhead costs totaled $800,000, of which $294,700 is fixed overhead.

Required:

1. Calculate the fixed overhead spending and volume variances.

Fixed Overhead Spending Variance $
Fixed Overhead Volume Variance $

2. Calculate the variable overhead spending and efficiency variances.

Variable Overhead Spending Variance $
Variable Overhead Efficiency Variance $

3. Prepare the journal entries that reflect the following:

  1. Assignment of overhead to production
  2. Recognition of the incurrence of actual overhead
  3. Recognition of overhead variances
  4. Closing out overhead variances, assuming they are not material

Note: Close the variances with a debit balance first. For compound entries, if an amount box does not require an entry, leave it blank or enter "0".

a.
b.
c.
d.
Answer #1

ANSWER:

1.

Fixed Overload Spending Variance=Actual Fixed Overhead Expense-Budgeted Fixed Overhead

=$294,700-300,000

=5,300(Favorable)

Fixed Overload Volume Variance=(Std. Absorption rate * Actual units) - (Std. Absorption rate * Budgeted unit)

=(0.30*900,000)-(0.30*1,000,000)

=$270,000-$300,000

=30,000(Favorable)

2.

Variable overhead spending variable =Actual variable overhead-Actual labor hours*standard rate

=$505,300-(190,000*2)

=$125,300(Unfavorable)

Variable overhead volume variance =(SH*SR)-(AH*SR)

=(200,000*2)-(190,000*2)

=400,000-380,000

=20,000(Favorable)

---------------------------------------------

As HOMEWORKLIB rule we answered only Two Questions, Thank you.

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************

)